### v1

@@ 7.2.1 Carbothermal reduction architecture
The CR plant can achieve the lowest required excavation rate among the examined architectures (Fig. 18). Moreover, the CR plant can be the lightest architecture when the operational period is one year (Figs. 9(a), 11(a), and 12(a)). However, this architecture depends the most on the import of resources from Earth (Fig. 7), which can make it significantly heavier (and costly) when the operational period is longer (Fig.9(b)). Another drawback of this architecture is the relatively high power consumption (see e.g., Figs.10 and 18).
@@ 7.2.2 Water extraction architecture
As the Monte Carlo analyses reveals, the WE plant has the widest range in the required regolith excavation rate (Fig. 18). Although it can still be the lightest plant architecture for three years’ operation (see, e.g., Figs. 9(b) and 12(b)), it requires a lot more mass to improve the robustness of the system (Table 5). If the water content in the excavated regolith is indeed as high as 30 wtas reported in [38], the expected regolith excavation becomes smaller (Fig. 20) and the expected power consumption also becomes smaller than the other architectures (Fig. 20).
@@ 7.2.3 Parallel Hybrid architecture
The performance, indicated by the power consumption, and regolith excavation, of the Parallel Hybrid architecture turned out to be between that of the CR the WE architectures. The required landed mass for the Parallel Hybrid architecture is not the lightest. However, the mass difference between the Parallel Hybrid architecture and the CR or the WE architecture not always significant (see e.g., Fig. 11(b)), indicating that one can gather a lot more information that cannot be obtained from the CR or WE architecture by adding a little mass. Therefore, having a hybrid pilot plant to gather more information to help decision-makers make more informed decisions for a full-scale plant could be beneficial. This potential flexibility in the operation and deployment of a hybrid plant should be explored in the future.
@@ 7.2.4 Series Hybrid architecture
The obvious benefit of the Series Hybrid architecture is the small regolith excavation rate with a small variance regardless of the water ice content (Figs. 18 and 20). This merit comes with a cost as the required mass which is often the heaviest among the examined architectures. This mass cost can become even larger when we consider the technological challenge of performing the CR process in a PSR.
@@ 8 Conclusion
The uncertainty in water ice content largely affects the required excavation rate. This suggests the urgent need to gather more information on lunar water ice. The Series Hybrid architecture turned out to be a reasonable choice in terms of the regolith excavation rate showing its merit regardless of the water ice content. However, the Series Hybrid architecture becomes the heaviest architecture in many cases.
@@ Acknowledgement
The authors would like to thank Dr. Joshua N. Rasera and Luka Malone for their reviews and thoughtful suggestions for improvement.
@@ Appendix A Comparisons with existing mass and power estimation
This appendix compares the mass and power of each subsystem estimated from the proposed model with values reported in [7, 23]. Table 6 lists the mass and power of each subsystem estimated using the proposed model and reported in Ref. [7]. Linne et al. scaled their plant using three independent plants with each producing 3.5 t of oxygen. To compare with this model, we scaled a plant to produce 3.5 t of oxygen and multiplied the mass and power by three. There are large differences in the mass of the tanks. This is likely due to the different assumptions in the reactant recycle rate. In our model, only 91% of the reactant is assumed to be recycled, whereas such an assumption is not made in Ref. [7]. Ref. [7] also modeled material transportation and size sorting subsystems in more detail than this paper, which leads to a large mass and power difference in the materials processing and handling subsystem.
Table 6: Comparison of the Carbothermal Reduction Plant with Linne et al. [7]
@@
Element Est. value Ref. [23] Est. value Ref. [23]
CR reactor 448 177 89,828 43,290
Desulfurization - 78 - 20
@@
Recycle loop compressor 2 2 204 1,482
Heat exchanger 2 17 - -
Condenser/electrolyzer H2O tank - 28
System gas tank w/ compressor - 18 - 585
H2 tank 192 8 - -
CH4 tank 91 8 - -
Materials processing & handling 54 350 2 793
Electrolyzer subsystem 13 199 16,274 12,480
Excavator 223 260 633 -
@@
Table 7 lists the mass and power of each subsystem estimated using the proposed model and reported in Ref. [23]. To compare these values, the assumptions here are adopted from Ref. [23]. The water content used here is 5 wt%, and the operational availability is adjusted to 225 days per lunar year. The ratio of O2 and H2 here is 6:1; therefore, the plant is scaled to produce 13.3 t of LOX and 1.7 t of LH2. The mass difference in the water extractor is due to the water extraction method; our extractor model is based on a batch operation, whereas Kleinhenz and Paz assumed continuous water extraction [23]. The required number of excavators is one based on our model, which is also different from the two excavators’ operation in Ref. [23]. The mass of gas dryers is modeled based on the in-house developed hardware at Johnson Space Center in Ref. [23], which causes large mass difference from the model developed in this paper. There is another large difference in the total mass of water tankers due to the number of tankers. While Kleinhenz and Paz assumed only two tankers are required [23], our model outputs a lighter total mass when 10 tankers are used. This mass can be further lighter by increasing the maximum number of tankers; however, operating a large number of tankers could increase the complexity of the total system.
Table 7: Comparison of the direct water extraction plant with Kleinhenz and Paz [23].
Mass [kg]∗1 Power [kW]∗1
Element Est. value Ref. [23] Est. value Ref. [23]
Water extractor 533 335 25.01 20.75
Excavator 86 160 0.49 0.21
O2 liquefaction 169 153 4.08 4.09
O2 cryostorage 242 213 - -
H2 liquefaction 1933 1780 25.94 21.17
H2 cryostorage 282 299 - -
H2 & O2 dryers 1 52 - 0.06
PEM electrolysis 10 48 22.32 22.04
Water tank 264 123 - 0.20
Water tankers 3742 1868 2.63 PEL: 0.38
PSR: 0.53
$\ast 1$: All values do not include 30% growth to compare them with Ref. [23].

### v2

@@ 7.2.1 Carbothermal reduction architecture
The CR plant requires a slow excavation rate (Fig. 20) due to the high oxygen yield of the process. Moreover, the CR plant can be the lightest architecture when the operational period is one year (Figs. 11(a), 13(a), and 14(a)) and the least power-consuming architecture in comparison depending on the water content (see Figs. 15 and 20). However, this architecture depends the most on the import of resources from Earth (Fig. 8), which can make it significantly heavier (and costly) when the operational period is longer (Fig.10).
@@ 7.2.2 Water extraction architecture
As the Monte Carlo analyses reveal, the WE plant has the widest range in the required regolith excavation rate (Fig. 20), required power (Fig. 20), and landed mass (Fig. 23) While it can still be the lightest plant architecture for operations lasting two years or longer (Fig. 10), achieving greater system robustness requires significantly more mass (Table 6).
If the water content in the excavated regolith is indeed as high as 30 wtas reported by Li et al. [38], the expected regolith excavation becomes smaller (Fig. 22) and the expected power consumption becomes the smallest (Fig. 22). However, if the water content is low, this architecture demands substantially higher power (Fig. 15). Therefore, decision-makers must carefully evaluate this option and ensure greater certainty in water content before implementation.
@@ 7.2.3 Parallel hybrid architecture
The performance, indicated by the power consumption, and regolith excavation, of the PH architecture turned out to be between that of the CR and the WE architectures (see e.g., Figs. 12 and 14). The required landed mass for the PH architecture is not the lightest. However, the mass difference between the PH architecture and the CR or the WE architecture not always significant (see e.g., Fig. 11), indicating that one can gather a lot more information that cannot be obtained from the CR or WE architecture by adding a little mass. Therefore, having a hybrid pilot plant to gather more information to help decision-makers make more informed decisions for a full-scale plant could be beneficial. This potential flexibility in the operation and deployment of a hybrid plant should be explored in the future.
@@ 7.2.4 Series hybrid architecture
The obvious benefit of the SH architecture is the small regolith excavation rate with a small variance regardless of the water ice content (Figs. 20 and 22). This merit comes with a cost as the required mass which is often the heaviest among the examined architectures (see e.g., Fig. 23). This mass cost can become even larger when we consider the technological challenge of performing the CR process in a PSR.
Regarding power consumption, this architecture is expected to require relatively small power in comparison regardless of the water ice content (Figs. 15, 20, and 22). Therefore, for risk-averse decision-makers, this architecture would be advisable.
7.3 Limitations and direction for future work
In this paper, the performances of different lunar ISRU architectures are compared quantitatively. As Kiewiet et al. [22] discussed, however, qualitative aspects of each technology, such as technological maturity should also be considered as part of the criteria. The maturity of the technologies and the complexity of each design were considered as criteria by Taylor and Carrier [25]. Similar analysis to these studies should be conducted in the future to help decision-makers consider other perspectives.
Furthermore, there are risks inherent in each plant architecture. All considered architectures employ water electrolysis to produce oxygen and hydrogen. There exists a risk of potential suspension of the products when a technological malfunction occurs. The benefit of deploying multiple smaller-scale plants similar to the study by Linne et al. [7] should be explored as well.
For the SH architecture, rejecting a large heat load in a PSR can affect the environment in multiple ways. It could lead to the sublimation of icy water in the PSR. It also could affect the icy regolith underneath the processing plant. Without carefully assessing these risks, the SH architecture should not be deployed on the Moon.
Global optimizations of the entire design and operations are left for future work. Adjusting certain parameters, such as the size of the molten zone in the CR process, may lead to a more optimal design. Additionally, design parameters can affect the operations of each architecture, potentially impacting both average and peak power consumption.
This paper also does not explore the flexibility in the operation enabled by a hybrid plant. Deploying a hybrid plant can gather more data than a single technology plant, indicating the better decisions to be made next. This flexible operation will be discussed in the future.
For a more detailed ISRU system design, improving the fidelity of certain subsystem models is desirable. Incorporating a parametric sizing model for the methanation chamber and integrating various beneficiation technologies may provide decision-makers with additional insights into system design and operations.
Finally, the inclusion of power systems and consideration of different technologies, such as microwave heating of icy regolith, would help decision-makers to better understand the trade space between different architectures. A similar approach to this paper, subsystem-level modeling, and uncertainty analysis, can be taken for such an extension, as well.
@@ 8 Conclusion
The uncertainty in water ice content largely affects the required excavation rate as well as power consumption. This suggests the urgent need to gather more information on lunar water ice. The series hybrid architecture turned out to be a reasonable choice in terms of the regolith excavation rate showing its merit regardless of the water ice content. This architecture also consumes relatively low power among all considered architectures. However, the series hybrid architecture becomes the heaviest architecture in many cases.
@@ Acknowledgement
The authors would like to thank Joshua N. Rasera, Ph.D., Luka Malone, and three anonymous reviewers for their careful reading and thoughtful suggestions for improvement.
[…]
@@ Appendix A Comparisons with existing mass and power estimation
This appendix compares the mass and power of each subsystem estimated from the proposed model with values reported in two past studies [7, 23]. Table A.1 lists the mass and power of each subsystem estimated using the proposed model and reported in Linne et al. [7]. Linne et al. scaled their plant using three independent plants with each producing 3.5 t of oxygen. To compare with this model, we scaled a plant to produce 3.5 t of oxygen and multiplied the mass and power by three. There are large differences in material transporting mass. In this paper, we have only considered an auger-based particle size separator, while multiple longer regolith transportation systems are considered by Linne et al. The PEM stack, too, shows a large difference in mass between our estimation and theirs. Note that the model used by Linne et al. is unclear, and thus, further assessment of the PEM stack is required.
Table A.1: Comparison of the Carbothermal Reduction Plant with Linne et al. [7]
@@
Element Est. value Linne et al. [7] Est. value Linne et al. [7]
CR reactor 572 177 48,270 43,290
Desulfurization (not modeled) 78 (not modeled) 20
@@
Recycle loop compressor 1 2 204 1,482
Heat exchanger 5 17 - -
Condenser/electrolyzer H2O tank (not modeled) 28 - -
System gas tank w/ compressor (not modeled) 18 (not modeled) 450
H2 tank 10 8 - -
CH4 tank 3 8 - -
Hoppers 94 123 - -
Materials size-sorting and transporting 2 227 12 610
Electrolysis pump 3 6 128 450
PEM stuck 8 152 16,145 9,150
Dryers 1 4 - -
Electrolysis valves and lines (not modeled) 38 - -
Excavator 223 276 1,266 -
@@
Table A.2 lists the mass and power of each subsystem estimated using the proposed model and reported in Kleinhenz and Paz [23]. To compare these values, the assumptions here are adopted from this reference. For instance, the water content used here is 5 wt%, and the operational availability is adjusted to 225 days per lunar year. The ratio of O2 and H2 here is 6:1; therefore, the plant is scaled to produce 13.3 t of LOX and 1.7 t of LH2. The mass difference in the water extractor is due to the water extraction method; our extractor model is based on a batch operation, whereas Kleinhenz and Paz assumed continuous water extraction. The required number of excavators is one based on our model, which is also different from the two excavators’ operation proposed by Kleinhenz and Paz. The mass of gas dryers is modeled based on the in-house developed hardware at Johnson Space Center in their study, which causes large mass difference from the model developed in this paper. There is another large difference in the total mass of water tankers due to the number of tankers. While Kleinhenz and Paz assumed only two tankers are required, our model outputs a lighter total mass when 10 tankers are used. This mass can be further lighter by increasing the maximum number of tankers; however, operating a large number of tankers could increase the complexity of the total system.
Table A.2: Comparison of the direct water extraction plant with Kleinhenz and Paz [23].
Mass [kg]∗1 Power [W]∗1
Element Est. value Kleinhenz and Paz [23] Est. value Kleinhenz and Paz [23]
Water extractor 487 335 25,014 20,749
Excavator 86 160 487 211
O2 liquefaction 159 153 4,800 4,093
O2 cryostorage 727 213 - -
H2 liquefaction 1923 1780 25,515 21,171
H2 cryostorage 902 299 - -
H2 & O2 dryers 1 52 - 58
PEM electrolysis 7 48 22,192 22,042
Water tank 266 123 - 199
Water tankers 3742 1868 2,633 PEL: 375
PSR: 534
$\ast 1$: All values do not include 30% growth to compare them with Kleinhenz and Paz [23].